# Extracting Sentence Embeddings from Pretrained Transformer Models

### v1

@@ 3.5.2 Semantic textual similarity (STS) tasks
STS assesses the degree to which two sentences are semantically equivalent to each other. A single sample consists of two sentences and a score ranging from 0 for no meaning overlap to 5 for meaning equivalence. The semantic textual similarity shared task has been held annually since 2012 up to 2017 [218, 219, 220, 221, 222, 223] and formed STS12-STS17 datasets. Carefully selected 8 628 samples from these contests formed the STS benchmark [223]. Table 3 shows details of the STS datasets, including SICK-Relatedness [224] and STS-B test sets, which we also use. Similarly to the work of [145], we also add (STR) [225], a recent semantic relatedness dataset created by comparative annotations.
@@
Table 3: Dataset statistics for STS and classification tasks. N is the number of text samples, C is the number of clusters, L/S is the imbalance number defined as the size of the largest class divided by that of the smallest class (for STS tasks the two classes are binned to 1 point label value length ones from the highest and lowest sides), $\|V\|$ is vocabulary size, Len is the average number of tokens in each text sample, Alpha is the percent of tokens that are alphabetic. Statistics with plain word tokenization are marked with “W” and with bert-base-uncased model tokenizer as “B”. For MRPC, SICK-R/E, STS-B, and STS tasks, statistics are calculated for tokenized and then concatenated sentences in each pair.
@@
SUBJ 10 000 2 1.0 22 636 15 912 25 28 85 78
MPQA 10 606 2 2.2 6 239 6 248 3 3 97 88
SST2 train 67 349 2 1.3 14 816 11 570 9 11 88 78
@@
MRPC train 4 076 2 2.1 16 112 12 061 44 50 81 75
test 1 725 2 2.0 10 092 8 471 44 50 82 75
@@
We follow the STS evaluation settings from [143]. First, the evaluation is kept unsupervised by not applying any additional regressors: cosine similarity between embeddings in a pair is taken directly as a model score for similarity. To find the degree of correlation between the annotated and model-given labels, Spearman’s rank correlation is used because it measures the rankings instead of the actual scores. Finally, for annual STS challenges, we concatenate all subsets and report the general Spearman correlation for that year (referred in [143] as “all”). As we used the SentEval toolkit [194], we had to implement concatenation ourselves since it only had “mean” and “wmean” settings available.
@@ 3.5.3 Downstream classification tasks
It includes sentiment prediction from Stanford Sentiment Treebank dataset SST2 [112], movie reviews MR [226], and customer product reviews CR [227]. In SUBJ [228] binary subjectivity status is labeled for sentences from movie reviews and plot summaries and in MPQA [229, 230] phrase-level opinion polarity from news articles is predicted. The last is MRPC [231], Microsoft Research Paraphrase Corpus from parallel news sources for the paraphrase detection task.
@@
SCICITE [232] is a domain-specific classification task that assigns one of three intent labels (background information, method, result comparison) to sentences collected from scientific articles citing other articles. Meanwhile, the SICK-E [224] dataset has labels for sentence pairs as entailed, contradiction, or neutral.
@@
It includes SST5 [112], a 5-level sentiment analysis dataset, and TREC [233] comprising classification of 6 types of questions. We also evaluate regular semantic textual relatedness and similarity tasks SICK-R [224] and STS-B [223] with classification, by splitting the real-valued similarity targets into 5 discrete class labels. For example, a [0,5] score of 3.6 goes to class 3 with weight 0.4 and also to class 4 with weight 0.6.
@@ 3.5.4 Isotropy
For each text embedding method, we calculated IsoScore for Wikitext-2 dataset from [201]. We split the data set into individual sentences (now samples) and omitted texts with less than 10 characters.
@@ 3.5.5 Alignment and uniformity
As a distribution of positive pairs $p_{\text{pos}}$ we used STS-B task train split pairs with a similarity of 5.0 and for $p_{\text{data}}$ we used sentences from all pairs.
@@ 4 Results
Here we present the results of our various experiments outlined in Section 3.
@@ 4.1 Token aggregation and post-processing techniques
We see that both aggregation and post-processing methods have a great impact on STS and clustering performance. For semantic textual similarity tasks, the best model was improved from 62.3 to 71.6 average Spearman correlation, while for clustering the best model was improved from 59.2 to 64.8 average accuracy.
Table 4: Effect of different token aggregation and post-processing methods on multiple text embedding models. Both correlation and accuracy scores range form 0 to 100 (from the worst to the best).
@@
We considered many post-processing techniques for sentence vectors and show the performance of the most popular and best-performing ones in Table 5 separately. Indeed, almost every method somehow improves the average Spearman correlation for STS tasks. For all-but-the-top (abtt), we varied the number of components to remove up to a hundred but settled on removing only the first two, as it was slightly better than the rest. Although highly credited in the literature, abtt-2 still got smaller scores than the others. Plainly averaging the token representations, the highest scores for both STS and clustering tasks were achieved using quantile-uniform normalization. We would also like to mention the simple normalization of vectors to the unit length, which was often beneficial for clustering tasks; and whitening normalization, which was learned on Wikitext-2 dataset, and was favorable with the random embeddings model. We trained other post-processing methods on Wikitext-2 too, yet they resulted in similar or slightly smaller scores.
Table 5: Performance of post-processing with plain token averaging.
@@
This result is logical because similarity or clustering tasks are done with the resulting text representations directly, while the classification is a supervised task learned on top. In the first case, the most informative components of the representations must be present in the largest principal components (i.e., constitute most of the variance in the data) for high scores, while the supervised logistic regression classifier can learn to extract them from the small principal components or ill-shaped representations on its own.
@@
An interesting result, as seen in Table 4, is the comparison of simply averaged BERT tokens from many contexts (the Avg. model) and word vectors from a specially-trained word2vec-style model BERT2Static (the B2S model) using the same BERT contexts, as described in Section 3.2.4. If no token aggregation and post-processing techniques are used, B2S is of similar performance at STS tasks (with average Spearman correlation of 56.7 versus 56.4), worse at clustering (with average clustering accuracy of 53.8 versus 55.2) and comparable at classification tasks (with average B2S accuracy of 76.9 versus 77.3) to the Avg. model. As we can see, the task performance differences are very small.
Apart from training, the second main difference between the Avg. and B2S models is tokenization. Avg. model uses sub-word tokens, same as BERT (tokens of which it averages), while B2S models use a vocabulary of the full most frequent words, 20 times the size of the BERT’s. Due to this discrepancy, token aggregations for BERT and B2S models are applied differently, as idf statistics for both tokenizations are different; also B2S does not have the equivalent of biased tokens such as BERT. However, it is very straightforward for B2S to filter out (do not use during averaging) a portion of the most frequent words. During our experiments, we varied this number and found that removing the 100 most frequent words works best, which improved the average Spearman correlation of STS tasks from 56.7 to 60.6 and the average accuracy of clustering tasks from 53.8 to 55.3 compared to full B2S, both with plain token averaging and no post-processing applied.
@@ 4.1.2 BERT versus random embeddings
The most dramatic increase in performance due to the two techniques is observed for the model of random embeddings. For STS tasks, the average Spearman correlation rises from 50.3 to 66.4, while for clustering tasks it increases from 36.3 to 49.5. Although accuracy for the latter tasks is still very low, for STS it is just mere 3.4 points below what BERT managed to accomplish with both the techniques applied.
If we were to look at the detailed performance on individual tasks in Appendix A.1, we would find several where random embeddings with the help of the two shaping techniques score higher than also-shaped BERT representations. For STS14 (Table 8) it is 68.8 versus 68.3 Spearman correlation, for googleTS (Table 16), stackoverflow (Table 18), and tweet (Table 19) datasets the clustering accuracy is correspondingly 69.5 vs 68.5, 70.6 vs 59.9, and 58.5 vs 55.1. This may indicate a smaller complexity of these particular datasets, where the task can largely be solved based on a small set of keywords. Yet it is important to note that it is achieved only with the help of aggregation and post-processing methods on top of the random embeddings.
The largest clustering accuracy difference between BERT and random embeddings is for agnews and searchsnippets datasets, with the best scores of 86.8 versus 43.4 and 82.9 versus 36.6 respectively. We think that the observed performance gap may be related to the vocabulary and text sizes of the datasets. Stackoverflow and tweet datasets, with random embeddings ahead of the BERT, has the smallest vocabulary sizes of 7 332 and 8 091 (see Table 2), while for agnews and searchsnippets has the largest vocabulary of 16 140 and 16 334 unique tokens respectively. Because random embeddings rely only on distinctness of tokens, with the increased amount of them, the probability of having exactly the same keywords in two texts drops. BERT embeddings having non-identical but semantically similar tokens similar representations helps in this case.
Having shorter text lengths (average text sizes 12 and 11 for stackoverflow and tweet respectively, versus 26 and 24 for agnews and searchsnippets respectively) may also help the random embeddings, because they do not average away so easily, and there is less context to consider.
@@ 4.1.3 Isotropy
Many post-processing methods have previously been proposed to improve the isotropy of representations. It was argued that representations from the BERT model fall into the narrow cone and, therefore, are anisotropic. Thus, by raising isotropy one can improve regular task performance.
We calculate the IsoScore isotropy metric on the Wikitext-2 dataset after applying various post-processing and token aggregation methods for multiple models. We exclude whitening post-processing, as it always produces representations with the IsoScore at the maximum of 100%, just due to its working principle. Then we compare the isotropy score to the STS, clustering, and classification task performance by calculating the Pearson correlation coefficient. For semantic textual similarity tasks, results are shown in Figure 1 and for clustering tasks in Figure 2.
For each method, the baseline IsoScore is very low: always below 5%. After applying weighted token aggregation techniques or post-processing, it is always increased. The smallest improvements of up to 7-8% score are observed for templated models T0 and T4, while the highest, over 80% are for random embeddings. For other models, the IsoScore reaches around 17%.
Why is the IsoScore of random embeddings improved so much compared to the other models? The answer may be related to the inner workings of RE. Once we generate the random embeddings, these token-level representations have a maximum possible isotropy. RE model represents text sequences as averages of these vectors, thus isotropy going from the token level to the document is reduced. However, token aggregation and post-processing techniques mitigate this isotropy loss, allowing for document embeddings to regain most of it back from the token-level ones. Other models start with already low isotropy for token-level representations and thus have less space to improve it.
@@
Six out of eight models have a moderate correlation (more than 54%) between Isoscore and the average Spearman correlation of semantic textual similarity tasks (see Figure 1). For clustering (see Figure 2) it is less apparent, with only 4 models reaching moderate Pearson correlation. However, for both STS and clustering tasks, the best score for each model is always reached by some improvement of IsoScore, compared to the isotropy score of the plain averaging setup.
@@
In contrast, the classification does not improve with token aggregation and post-processing techniques, therefore, we do not observe correlation (and do not show it here).
@@ 4.1.4 Alignment and uniformity
We can observe in Figure 3 that the zscore post-processing always makes alignment worse, while normalization of embeddings almost always does the same for the uniformity of the representations. Excluding the random embeddings model, the best alignment and uniformity properties are almost always with plain averaging and no post-processing applied.
@@ 4.2 Using prompts
The normal use of prompts, as presented in [15], is to place the sentence [X] inside a template as This sentence : "[X]" means [MASK] and use only the vector of [MASK] token from the final layer as the full representation of [X]. Our experiments show that in general using prompts is beneficial, however, we found some ways to improve performance even more.
@@
The use of the additional text template around the text enriches the target text representation. Let us compare the performance of 12th layer averaged token representations with and without a template (see Figure 4, T0 avg. and T4 avg. with a template versus BERT avg. without). For all 3 groups of tasks – semantic textual similarity, clustering, and classification – the T0 template achieves 61.3 average Spearman correlation, 53.5 clustering accuracy, and 80.5 classification accuracy; the T4 template reaches 62.3, 55.0, and 79.9, while a regular, non-templated text gets just 53.2, 50.6, and 79.0 correspondingly. This shows that the use of the templates allows the model to enrich target text representation. However, this effect peaks at the last 12th layer and the achieved performance is similar to the first+last layer combination of the regular non-templated vectors.
Performance of using only the [MASK] token also peaks at the 12th layer, so we investigated what influence it has in the enrichment of templated target text representation. Could it be that this token is the main culprit for better performance of averaged templated text representations? To answer this question we tried to omit the [MASK] token from averaging (see Figure 4 T0 avg. no [MASK] and T4 avg. no [MASK]). For the T4 template and 12th layer representations, averaging all tokens yields 62.3 average Spearman correlation, 55.0 clustering accuracy, and 79.9 classification accuracy; dropping the [MASK] token from averaging: 58.7, 53.3, and 79.8; while non-templated performance is 53.2, 50.6, and 79.0 correspondingly. We see that omitting the [MASK] token in averaging indeed hurts the performance. On the other hand, the results show that it is responsible only for approximately half of the improvements, with the other half coming from the other tokens in the templated text.
One good reason to use averaging instead of only the [MASK] token is that then token weighting can be also applied. As we already showed in Table 4), the T4 template together with $idf_{t}^{T}$ token weighting and uniform quantile post-processing allowed to reach the average Spearman correlation of 71.6, which was the best among the tried methods. Also, we observed that post-processing on text representations from [MASK] token was not as effective as from averaged ones. This also suggests that all tokens in templated texts have richer representations.
Figure 4: Layer-wise performance of templated models T0 and T4, as well as BERT versus RE with no weighting or post-processing. The average performance of STS, clustering, and classification tasks is shown by the lines while shadow areas correspond to the standard deviation. We also show first+last aggregation over layers as the last tick $1\atop 12$ on the horizontal axis.
@@
One of the ways we sought to improve BERT transformer model representations is to combine embeddings of the regular BERT with the ones averaged over multiple contexts. That is, for each token in each BERT layer, we collected many different contexts, and the averaged vector became the vector of the Avg. model. We then combined BERT and Avg. model, according to the parameter $w$ that shows the fraction of Avg. model representations in the resulting vector $v$:
@@
We have varied the $w$ parameter to also the negative values in (10) to see if subtracting (instead of adding) the context-average representations from the context-aware ones helps.
@@
Figure 5: BERT+Avg. model performance dependence on the weight $w$ of Avg. model and layer, from which (for both models) representations are used. To the right of the black line on the horizontal axis average aggregation of multiple layers is also shown. Tokens are simply averaged and no post-processing is used. The horizontal line with $w=0.0$ corresponds to a regular Bert (B) model, $w=0.5$ is B+Avg., and $w=1.0$ is the Avg. model. The white $\times$ marks the maximum value.
@@
Meanwhile, as we see in the same table, the average accuracy of classification tasks without additional techniques applied to the combination of the first and last layers is similar between only BERT (79.9%) and BERT+Avg. (79.7%). However, as we see in Figure 5, differently from the first+last optimal layer setting for BERT, BERT+Avg. has a sweet spot in the 10th layer with an average classification accuracy of 80.5%, surpassing the one of BERT by 0.6%.
@@
As we already mentioned, STS tasks preferred the regular BERT ($w=0$) to the BERT+Avg. on average. However, for several individual semantic textual similarity datasets, the best weights turned out to be even negative. This is the case for STS15 ($w=-0.25$), and STS16 ($w=-0.75$). The same negative weight of $w=-0.5$ preference was also observed for the searchsnippets clustering dataset. Although the Spearman correlation for three STS tasks was only up to several percent higher, for searchsnippets the clustering accuracy increased to a staggering 80.2% from 72.2% of $w=0.0$. This shows that the determined values of $w=0.0$ for STS and $w=0.5$ for clustering and classification are not universal and for a small percentage of tasks can differ. For more BERT+Avg. task-wise details, see Appendix A.2 figures for STS (Figure 6), clustering (Figure 7), and classification (Figure 8) performance.

### v2

@@ 3.5.2 Semantic Textual Similarity (STS) Tasks
STS assesses the degree to which two sentences are semantically equivalent to each other. A single sample consists of two sentences and a score ranging from 0 for no meaning overlap to 5 for meaning equivalence. The semantic textual similarity shared task has been held annually since 2012 up to 2017 [218, 219, 220, 221, 222, 223] and formed STS12-STS17 datasets. A total of 8628 carefully collected samples from these contests formed the STS benchmark [223]. Table 3 shows details of the STS datasets, including SICK-Relatedness [224] and STS-B test sets, which we also use. Similarly to [145], we also add (STR) [225], a recent semantic relatedness dataset created by comparative annotations.
@@
Table 3: Dataset statistics for STS and classification tasks. N is the number of text samples, C is the number of clusters, L/S is the imbalance number, defined as the size of the largest class divided by that of the smallest class (for STS tasks, the two classes are binned to 1 point label value length ones from the highest and lowest sides), $\|V\|$ is vocabulary size, Len is the average number of tokens in each text sample, Alpha is the percent of tokens that are alphabetic. Statistics with plain word tokenization are marked with “W” and with bert-base-uncased model tokenizer as “B”. For MRPC, SICK-R/E, STS-B, and STS tasks, statistics are calculated for tokenized and then concatenated sentences in each pair.
@@
SUBJ 10,000 2 1.0 22,636 15,912 25 28 85 78
MPQA 10,606 2 2.2 6239 6248 3 3 97 88
SST2 train 67,349 2 1.3 14,816 11,570 9 11 88 78
@@
MRPC train 4076 2 2.1 16,112 12,061 44 50 81 75
test 1725 2 2.0 10,092 8471 44 50 82 75
@@
We follow the STS evaluation settings from [143]. First, the evaluation is kept unsupervised by not applying any additional regressors; cosine similarity between embeddings in a pair is taken directly as a model score for similarity. To find the degree of correlation between the annotated and model-given labels, Spearman’s rank correlation is used because it measures the rankings instead of the actual scores. Finally, for annual STS challenges, we concatenate all subsets and report the general Spearman correlation for that year (referred to in [143] as “all”). As we used the SentEval toolkit [194], we had to implement concatenation ourselves since it only had “mean” and “wmean” settings available.
@@ 3.5.3 Downstream Classification Tasks
It includes sentiment prediction from Stanford Sentiment Treebank dataset SST2 [112], movie reviews MR [226], and customer product reviews CR [227]. In SUBJ [228], binary subjectivity status is labeled for sentences from movie reviews and plot summaries, and in MPQA [229, 230], phrase-level opinion polarity from news articles is predicted. The last is MRPC [231], the Microsoft Research Paraphrase Corpus, from parallel news sources for the paraphrase detection task.
@@
SCICITE [232] is a domain-specific classification task that assigns one of three intent labels (“background information”, “method”, “result comparison”) to sentences collected from scientific articles citing other articles. Meanwhile, the SICK-E [224] dataset has labels for sentence pairs as “entailed”, “contradiction”, or “neutral”.
@@
It includes SST5 [112], a 5-level sentiment analysis dataset, and TREC [233], comprising classification of 6 types of questions. We also evaluate regular semantic textual relatedness and similarity tasks SICK-R [224] and STS-B [223] with classification, by splitting the real-valued similarity targets into 5 discrete class labels. For example, a [0,5] score of 3.6 goes to class 3 with weight 0.4 and also to class 4 with weight 0.6.
@@ 3.5.4 Isotropy
For each sentence embedding method, we calculated IsoScore for the Wikitext-2 dataset from [201]. We split the dataset into individual sentences (now samples) and omitted texts with less than 10 characters.
@@ 3.5.5 Alignment and Uniformity
As a distribution of positive pairs $p_{\text{pos}}$, we used STS-B task training split pairs with a similarity of 5.0, and for $p_{\text{data}}$, we used sentences from all pairs.
@@ 4 Results
Here, we present the results of our various experiments outlined in Section 3.
@@ 4.1 Token Aggregation and Post-Processing Techniques
We see that both aggregation and post-processing methods have a great impact on STS and clustering performance. For semantic textual similarity tasks, the best model was improved from 62.3 to 71.6 average Spearman correlation, while for clustering, the best model was improved from 59.2 to 64.8 average accuracy.
Table 4: Effect of different token aggregation and post-processing methods on multiple text embedding models. Both correlation and accuracy scores range from 0 to 100 (from the worst to the best). We also added results for SimCSE-BERT model [143], which is fine-tuned on NLI data supervision. The best result for each model is bolded, while underlined results are the best across all the models.
@@
SimCSE performance: 59.8 and 64.0 (avg. + quantile-uW, first + last layers)
@@
SimCSE performance: 86.5
Our techniques can even improve the dedicated SimCSE model [143], which was fine-tuned on NLI data. Its main strength lies in semantic textual similarity tasks, where it leads with over 10% difference. However, for clustering tasks, its average accuracy is similar to the other evaluated models at 59.8% and improves up to 64.0%, if we apply the best-performing techniques. This showcases a general tendency that the top-performing models are very good only in a narrow subset of tasks and highlights the importance of our more general methods.
@@
We considered many post-processing techniques for sentence vectors and show the performance of the most popular and best-performing ones in Table 5 separately. Indeed, almost every method somehow improves the average Spearman correlation for STS tasks. For all-but-the-top (abtt), we varied the number of components to remove up to a hundred but settled on removing only the first two, as it was slightly better than the rest. Although highly credited in the literature, abtt-2 still obtained smaller scores than the others. Plainly averaging the token representations, the highest scores for both STS and clustering tasks were achieved using quantile-uniform normalization. We would also like to mention the simple normalization of vectors to the unit length, which was often beneficial for clustering tasks, and whitening normalization, which was learned on Wikitext-2 dataset, and was favorable with the random embeddings model. We trained other post-processing methods on Wikitext-2 too, yet they resulted in similar or slightly smaller scores.
Table 5: Performance of post-processing with plain token averaging. abtt-2 is the ABTT method with top 2 principal components removed. The best result for each model is bolded, while underlined results are the best across all the models.
@@
This result is logical because similarity or clustering tasks are carried out with the resulting text representations directly, while the classification is a supervised task learned on top. In the first case, the most informative components of the representations must be present in the largest principal components (i.e., constitute most of the variance in the data) for high scores, while the supervised logistic regression classifier can learn to extract them from the small principal components or ill-shaped representations on its own.
@@
An interesting result, as seen in Table 4, is the comparison of simply averaged BERT tokens from many contexts (the Avg. model) and word vectors from a specially trained Word2Vec-style model BERT2Static (the B2S model) using the same BERT contexts, as described in Section 3.2.4. If no token aggregation and post-processing techniques are used, B2S is of similar performance in STS tasks (with average Spearman correlation of 56.7 versus 56.4), worse at clustering (with average clustering accuracy of 53.8 versus 55.2), and comparable in classification tasks (with average B2S accuracy of 76.9 versus 77.3) to the Avg. model. As we can see, the task performance differences are very small.
Apart from training, the second main difference between the Avg. and B2S models is tokenization. The Avg. model uses sub-word tokens, the same as BERT (tokens of which it averages), while B2S models use a vocabulary of the full set of the most frequent words, 20 times the size of the BERT’s. Due to this discrepancy, token aggregations for BERT and B2S models are applied differently, as idf statistics for both tokenizations are different; also, B2S does not have the equivalence of biased tokens such as BERT. However, it is very straightforward for B2S to filter out (do not use during averaging) a portion of the most frequent words. During our experiments, we varied this number and found that removing the 100 most frequent words works best, which improved the average Spearman correlation of STS tasks from 56.7 to 60.6 and the average accuracy of clustering tasks from 53.8 to 55.3 compared to full B2S, both with plain token averaging and no post-processing applied.
@@ 4.1.2 BERT versus Random Embeddings
The most dramatic increase in performance due to the two techniques is observed for the model of random embeddings. For STS tasks, the average Spearman correlation rises from 50.3 to 66.4, while for clustering tasks, it increases from 36.3 to 49.5. Although accuracy for the latter tasks is still very low, for STS, it is just a mere 3.4 points below what BERT managed to accomplish with both techniques applied.
If we were to look at the detailed performance on individual tasks in Appendix A.1, we would find several where random embeddings with the help of the two shaping techniques score higher than the also-shaped BERT representations. For STS14 (Table 8), it is 68.8 versus 68.3 Spearman correlation, while for the googleTS (Table 16), stackoverflow (Table 18), and tweet (Table 19) datasets, the clustering accuracy is, respectively, 69.5 vs. 68.5, 70.6 vs. 59.9, and 58.5 vs. 55.1. This may indicate a smaller complexity of these particular datasets, where the task can largely be solved based on a small set of keywords. Also, if the texts do not contain the natural language of the type the BERT was pre-trained on (e.g., they contain code), the model cannot properly contextualize the tokens, and the random embeddings without contexts work better. Yet it is important to note that it is achieved only with the help of aggregation and post-processing methods on top of the random embeddings.
The largest clustering accuracy difference between BERT and random embeddings is for the agnews and searchsnippets datasets, with the best scores of 86.8 versus 43.4 and 82.9 versus 36.6, respectively. We think that the observed performance gap may be related to the vocabulary and text sizes of the datasets. Stackoverflow and tweet datasets, with random embeddings ahead of the BERT, have the smallest vocabulary sizes of 7332 and 8091 (see Table 2), while agnews and searchsnippets have the largest vocabulary sizes of 16,140 and 16,334 unique tokens, respectively. Because random embeddings rely only on distinctness of tokens, with the increased amount of them, the probability of having exactly the same keywords in two texts drops. BERT embeddings having non-identical but semantically similar tokens and similar representations helps in this case.
Having shorter text lengths (average text sizes of 12 and 11 for stackoverflow and tweet, respectively, versus 26 and 24 for agnews and searchsnippets, respectively) may also help the random embeddings, because they do not average away so easily, and there is less context to consider.
@@ 4.1.3 Isotropy
Many post-processing methods have previously been proposed to improve the isotropy of representations. It was argued that representations from the BERT model fall into the narrow cone and, therefore, are anisotropic. Thus, by raising isotropy, one can improve regular task performance.
We calculate the IsoScore isotropy metric on the Wikitext-2 dataset after applying various post-processing and token aggregation methods for multiple models. We exclude whitening post-processing, as it always produces representations with the IsoScore at the maximum of 100%, just due to its working principle. Then, we compare the isotropy score to the STS, clustering, and classification task performance by calculating the Pearson correlation coefficient. For semantic textual similarity tasks, results are shown in Figure 1, and for clustering tasks, results are shown in Figure 2.
For each method, the baseline IsoScore is very low: always below 5%. After applying weighted token aggregation techniques or post-processing, it is always increased. The smallest improvements of up to 7-8% score are observed for templated models T0 and T4, while the highest, over 80%, are for random embeddings. For other models, the IsoScore reaches around 17%.
Why is the IsoScore of random embeddings improved so much compared to the other models? The answer may be related to the inner workings of RE. Once we generate the random embeddings, these token-level representations have a maximum possible isotropy. The RE model represents text sequences as averages of these vectors; thus, isotropy going from the token level to the document is reduced. However, token aggregation and post-processing techniques mitigate this isotropy loss, allowing for document embeddings to regain most of it back from the token-level ones. Other models start with already low isotropy for token-level representations and thus have less space to improve it.
@@
Six out of eight models have a moderate correlation (more than 54%) between Isoscore and the average Spearman correlation of semantic textual similarity tasks (see Figure 1). For clustering (see Figure 2), it is less apparent, with only four models reaching moderate Pearson correlation. However, for both STS and clustering tasks, the best score for each model is always reached by some improvement of IsoScore, compared to the isotropy score of the plain averaging setup.
@@
In contrast, the classification does not improve with token aggregation and post-processing techniques; therefore, we do not observe correlation (and do not show it here).
@@ 4.1.4 Alignment and Uniformity
We can observe in Figure 3 that the z-score post-processing always makes alignment worse, while normalization of embeddings almost always does the same for the uniformity of the representations. Excluding the random embeddings model, the best alignment and uniformity properties are almost always with plain averaging and no post-processing applied.
@@ 4.2 Using Prompts
The normal use of prompts, as presented in [15], is to place the sentence [X] inside a template as “This sentence: "[X]" means [MASK]” and use only the vector of [MASK] token from the final layer as the full representation of [X]. Our experiments show that in general, using prompts is beneficial; however, we found some ways to improve performance even more.
@@
The use of the additional text template around the text enriches the target text representation. Let us compare the performance of 12th layer averaged token representations with and without a template (see Figure 4, T0 avg. and T4 avg. with a template versus BERT avg. without). For all 3 groups of tasks—semantic textual similarity, clustering, and classification—the T0 template achieves 61.3 average Spearman correlation, 53.5 clustering accuracy, and 80.5 classification accuracy; the T4 template reaches 62.3, 55.0, and 79.9, while a regular, non-templated text obtains just 53.2, 50.6, and 79.0, respectively. This shows that the use of the templates allows the model to enrich target text representation. However, this effect peaks at the last, 12th layer, and the achieved performance is similar to the first + last layer combination of the regular non-templated vectors.
Performance of using only the [MASK] token also peaks at the 12th layer, so we investigated what influence it has in the enrichment of templated target text representation. Could it be that this token is the main culprit for better performance of averaged templated text representations? To answer this question, we tried to omit the [MASK] token from averaging (see Figure 4 T0 avg. no [MASK] and T4 avg. no [MASK]). For the T4 template and 12th layer representations, averaging all tokens yields 62.3 average Spearman correlation, 55.0 clustering accuracy, and 79.9 classification accuracy; dropping the [MASK] token from averaging yields 58.7, 53.3, and 79.8; non-templated performance is 53.2, 50.6, and 79.0, respectively. We see that omitting the [MASK] token in averaging indeed hurts the performance. On the other hand, the results show that it is responsible only for approximately half of the improvements, with the other half coming from the other tokens in the templated text.
One good reason to use averaging instead of only the [MASK] token is that then token weighting can be also applied. As we already showed in Table 4), the T4 template together with $idf_{t}^{T}$ token weighting and uniform quantile post-processing allowed us to reach the average Spearman correlation of 71.6, which was the best among the tried methods. Also, we observed that post-processing on text representations from [MASK] token was not as effective as from averaged ones. This also suggests that all tokens in templated texts have richer representations.
Figure 4: Layer-wise performance of templated models T0 (subfigures a, d, g) and T4 (b, e, h), as well as BERT versus RE with no weighting or post-processing (c, f, i). The average performance of STS (a, b, c), clustering (d, e, f), and classification tasks (g, h, i) is shown by the lines, while shadow areas correspond to the standard deviation. We also show first + last aggregation over layers as the last tick $1\atop 12$ on the horizontal axis.
@@
One of the ways we sought to improve BERT transformer model representations is to combine embeddings of the regular BERT with the ones averaged over multiple contexts. That is, for each token in each BERT layer, we collected many different contexts, and the averaged vector became the vector of the Avg. model. We then combined BERT and Avg. model, according to the parameter $w$, which shows the fraction of Avg. model representations in the resulting vector $v$:
@@
We have also varied the $w$ parameter to the negative values in (10) to see if subtracting (instead of adding) the context-average representations from the context-aware ones helps.
@@
Figure 5: BERT + Avg. model performance dependence on the weight $w$ of Avg. model and layer, from which (for both models) representations are used. To the right of the black line on the horizontal axis, average aggregation of multiple layers is also shown. Tokens are simply averaged and no post-processing is used. The horizontal line with $w=0.0$ corresponds to a regular Bert (B) model, $w=0.5$ is B + Avg., and $w=1.0$ is the Avg. model. The white $\times$ marks the maximum value.
@@
Meanwhile, as we see in the same table, the average accuracy of classification tasks without additional techniques applied to the combination of the first and last layers is similar between only BERT (79.9%) and BERT + Avg. (79.7%). However, as we see in Figure 5, differently from the first + last optimal layer setting for BERT, BERT + Avg. has a sweet spot in the 10th layer with an average classification accuracy of 80.5%, surpassing that of BERT by 0.6%.
@@
As we already mentioned, STS tasks preferred the regular BERT ($w=0$) to the BERT + Avg. on average. However, for several individual semantic textual similarity datasets, the best weights turned out to be even negative. This is the case for STS15 ($w=-0.25$) and STS16 ($w=-0.75$). The same negative weight of $w=-0.5$ preference was also observed for the searchsnippets clustering dataset. Although the Spearman correlation for three STS tasks was only up to several percent higher, for searchsnippets, the clustering accuracy increased to a staggering 80.2% from 72.2% of $w=0.0$. This shows that the determined values of $w=0.0$ for STS and $w=0.5$ for clustering and classification are not universal, and for a small percentage of tasks, they can differ. For more BERT + Avg. task-wise details, see Appendix A.2 figures for STS (Figure 6), clustering (Figure 7), and classification (Figure 8) performance.